# Fine-tuning Large Language Models for Entity Matching

### v1

@@ Abstract.
Generative large language models (LLMs) are a promising alternative to pre-trained language models for entity matching due to their high zero-shot performance and their ability to generalize to unseen entities. Existing research on using LLMs for entity matching has focused on prompt engineering and in-context learning. This paper explores the potential of fine-tuning LLMs for entity matching. We analyze fine-tuning along two dimensions: 1) The representation of training examples, where we experiment with adding different types of LLM-generated explanations to the training set, and 2) the selection and generation of training examples using LLMs. In addition to the matching performance on the source dataset, we investigate how fine-tuning affects the model’s ability to generalize to other in-domain datasets as well as across topical domains. Our experiments show that fine-tuning significantly improves the performance of the smaller models while the results for the larger models are mixed. Fine-tuning also improves the generalization to in-domain datasets while hurting cross-domain transfer. We show that adding structured explanations to the training set has a positive impact on the performance of three out of four LLMs, while the proposed example selection and generation methods only improve the performance of Llama 3.1 8B while decreasing the performance of GPT-4o Mini.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
Artifact Availability:
The source code, data, and/or other artifacts have been made available at %****␣main.tex␣Line␣25␣****%leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/wbsg-uni-mannheim/TailorMatch.
1. Introduction
@@
Entity matching (Barlaug and Gulla, 2021; Christen, 2012; Elmagarmid et al., 2007) is the task of identifying entity descriptions in different data sources that refer to the same real-world entity. Entity matching is a central step in data integration pipelines (Christophides et al., 2020). Many state-of-the-art entity matching methods build on pre-trained language models (PLMs) (Vaswani et al., 2017), such as BERT or RoBERTa (Li et al., 2020; Peeters and Bizer, 2021; Peeters et al., 2024; Zeakis et al., 2023). Recent work on using generative large language models (LLMs) (Zhao et al., 2023), such as GPT, Llama, Gemini, or Mixtral, for entity matching has shown that LLMs have higher zero-shot performance compared to PLMs and generalize better to unseen entities (Narayan et al., 2022a; Peeters and Bizer, 2023, 2024). Most research on using LLMs for entity matching focuses on prompt engineering and in-context learning (Narayan et al., 2022a; Peeters and Bizer, 2023; Fan et al., 2023). Initial papers have started to investigate the potential of improving beyond in-context learning by fine-tuning LLMs for entity matching (Peeters and Bizer, 2024; Zhang et al., 2023). This paper builds on this work and presents a deeper investigation of the utility of fine-tuning LLMs for entity matching. Besides the effect of fine-tuning on the matching performance, we investigate how fine-tuning influences the model’s ability to generalize to different in-domain datasets as well as across domains. Our investigation focuses on two dimensions: First, the representation of training examples, where we investigate the effect of augmenting each training example with textual as well as structured explanations about why a pair of entity descriptions matches or not. The second dimension is the selection and generation of examples for the training set that is used for fine-tuning. We experiment with using LLMs to filter misleading examples from the training set as well as to generate additional training examples. We further investigate the error-based selection of training examples. Figure 1 gives an overview of the fine-tuning and inference setup. The effects of fine-tuning are compared between open-source (Llama-3.1) and proprietary (GPT-4o) models. While we test models of varying sizes, the focus of the investigation lies on smaller models, as these models are less resource-intensive to deploy.
@@
Figure 1. Fine-tuning and inference setup including explanation generation, example generation, and filtration.
@@
  2. (2)
    In-domain and Cross-domain Generalization: We assess the effectiveness of fine-tuned models in both in-domain and cross-domain settings. The experiments show that fine-tuned models often generalize to in-domain datasets, while their performance in cross-domain settings falls short of the zero-shot baseline in the majority of cases.
  3. (3)
@@
  4. (4)
    Example Selection and Generation: We introduce different approaches for using LLMs for filtering the training set and for the generation of new training examples. Both techniques improve the performance after fine-tuning, with example generation particularly enhancing in-domain generalization.
  5. (5)
    Prompt Sensitivity: We examine how fine-tuning affects the model’s sensitivity to variations of the prompt. Our experiments show that fine-tuning generally reduces prompt sensitivity, with the introduction of structured explanations further stabilizing model performance across different prompts.
2. Experimental Setup
This section provides details on the models, benchmark datasets, fine-tuning setup, and the performance evaluation that we use for the experiments. All artefacts are provided in the project repository, meaning that all experiments can be replicated.
@@ 2. Experimental Setup
  1. (1)
    gpt-4o-2024-08-06: Released in May 2024, this model is OpenAI’s flagship LLM and currently used by ChatGPT.
  2. (2)
    gpt-4o-mini-2024-07-18: This hosted LLM, released in July 2024, has lower usage fees compared to GPT-4o. Although the exact model size is not disclosed, the naming convention indicates that it is smaller than GPT-4o.
  3. (3)
@@
  4. (4)
    Meta-Llama-3.1-70B-Instruct222meta-llama/Meta-Llama-3.1-70B-Instruct: Also developed by Meta, this model is medium-sized within the Llama series. Further details are available in the accompanying paper by Meta (Dubey et al., 2024).
We use the OpenAI batch API [^3] https://platform.openai.com/docs/guides/batch and Hugging Face’s Transformers [^4] https://huggingface.co/docs/transformers/en/index library to prompt both hosted and local models. In both cases, the generation temperature is set at zero to minimize randomness. The open-source models run on local machines with varying hardware configurations, including an AMD EPYC 7413 processor, 1024GB RAM, and up to eight Nvidia RTX6000 GPUs.
Datasets: We evaluate in-domain and cross-domain generalization using benchmark datasets from two topical domains (Köpcke et al., 2010; Peeters et al., 2024): products and scholarly works. The datasets are selected based on three criteria: First, we chose datasets with at least 150 matches in the test set (see Table 1) to ensure the stability of the performance measurement. Second, we prioritize datasets that contain a significant number of challenging corner case record pairs. These corner cases are either matching or non-matching pairs that closely resemble the opposite class due to very similar or dissimilar surface forms (Peeters et al., 2024). Finally, to evaluate cross-domain generalization, we chose datasets from two topical domains. We use the following datasets:
@@
  • Walmart-Amazon: The Walmart-Amazon benchmark includes products from both Walmart and Amazon, with categories comparable to those in WDC Products and Abt-Buy.
@@
Table 1. Dataset statistics for the training, validation, and test sets.
@@
Fine-tuning Setup: In order to keep computational costs manageable, we use the default hyperparameter setting recommended by the model providers and do not perform a hyperparameter search for each experiment. The fine-tuning setup differs for hosted and open-source models. We use the following default hyperparameter settings for the hosted OpenAI models: learning rate multiplier 1.8, batch size 16, and a constant random seed for reproducibility. We use the following fine-tuning setup for the open-source models: We employed Low-Rank Adaptation (LoRA) with an alpha of 16 to allow moderate adjustments to model weights, a dropout rate of 0.1 to prevent overfitting, and a rank (r) of 64 to balance performance and computational efficiency. The learning rate was configured at 2e-4. To ensure reproducibility, a consistent random seed was used across all libraries, and CUDA was set to deterministic mode. Both open-source and hosted models are trained for 10 epochs, with checkpoints generated after each epoch. Open-source models are validated at each checkpoint using custom evaluation callbacks. In contrast, OpenAI models provided only two intermediate checkpoints plus the final version, limiting the validation process.
Evaluation: We use the standard evaluation metrics precision, recall, and F1-score, following related work (Barlaug and Gulla, 2021). All subsequent tables show the F1 performance of a model in the given setting. The precision and recall results of all experiments are provided in the repository accompanying this paper. Model responses in natural language are evaluated using Narayan et al.’s (Narayan et al., 2022a) method of parsing responses to contain “yes” or “no”.
3. Standard Fine-Tuning
In order to establish baselines, this section reports the results of a series of fine-tuning experiments using a simple, straight-forward representation of the training examples. Figure 2 shows two training examples from this fine-tuning set. Each example is represented by a prompt inquiring whether the entity descriptions refer to the same entity, followed by the two descriptions and the completion “Yes” or “No”, depending on whether the descriptions are a match or a non-match.
Figure 2. Representation of two training examples in the standard fine-tuning setup.
Table 2 reports the results of the standard fine-tuning experiments. Each row reports the results for a specific LLM (column Model) which was fine-tuned using a specific training set (column Training set). The following columns contain the results of evaluating this fine-tuned model on different test sets. The “zero-shot” rows report results for the specific models without any fine-tuning.
Table 2. F1 scores after standard fine-tuning. The rows determine the model/training set combination, while the columns contain results for different test sets. The best result of each model per test set is set bold, the second best result is underlined.
Model Training set Product Domain Transfer Scholar Domain Transfer
@@ 3.1. Effectiveness
Smaller Models: Llama-8B shows an average F1 score gain of 17.31 points over the zero-shot baseline (Table 2, upper section). Five out of six fine-tuned Llama 8B models significantly improve on their zero-shot baselines, with Abt-Buy achieving the highest gain of 30.77 points (Table 2, A-B column). The Amazon-Google model is an outlier, only exceeding the baseline by 0.84 points, underperforming compared to other product domain models (Table 2, A-G column). GPT-4o-mini shows an average F1 score increase of 11.72 points, a significant gain given the models high zero-shot performance (Table 2, , rows gpt-4o-m). Unlike Llama-8B, GPT-4o-mini demonstrates significant performance gains on the Amazon-Google dataset, achieving the highest absolute performance improvement among all 4o-mini models as presented in the A-G column of Table 2.
@@
Large Models: Given resource constraints, we fine-tune the larger models only on the WDC small training set. For the Llama 70B model fine-tuning results in a decrease in performance, with a reduction of 2.53 points in F1 score on the WDC test set, as shown in the WDC column of Table 2. In contrast, GPT-4o, which starts with a similar zero-shot score as GPT-4o-mini, continues the trend of fine-tuning improving performance. It achieves a 5.43-point increase to a F1 score of 87.07 on the WDC dataset, the highest result so far.
3.2. Generalization
@@
We evaluate two types of generalization: in-domain, where performance is assessed across datasets within the same topical domain, and cross-domain, where a model is trained using a training set from one domain and tested with a test set from a different topical domain (e.g., from products to scholar).
Transfer Gain: We quantify the capability of a fine-tuned model to generalize to target datasets by relating the performance gains of the fine-tuned model on the target datasets to the performance gain that is achieved by fine-tuning specialized models with training examples from each target dataset. The resulting transfer gain is calculated by dividing the average performance gain on target datasets by the average gain achieved by models which were fine-tuned specifically on those datasets.
[…]
For example, the generalization gain of the WDC model is 72% (Table 2, row Llama 8B/WDC, column Transfer Gain) and is calculated by dividing the average gain of the WDC model of 10.52 points on the product datasets by the average gain of specialized models for these datasets, which is 18.41 points F1.
In-domain Generalization: In the product domain, five out of six Llama 8B models improve performance on both their own and other datasets. Even the underperforming Amazon-Google model shows a gain on the Abt-Buy dataset. However, it is the only model to experience a decline in performance in an in-domain generalization setting, as indicated by the WDC column in Table 2. On average, Llama 8B models that were trained on a source dataset achieve 59% of the performance in an in-domain transfer scenario compared to models that were fine-tuned directly for the target datasets. GPT-4o-mini also exhibits some ability to transfer between different product datasets, but to a lesser extent, achieving only a 15% performance increase relative to dataset-specific models. This lower average is primarily influenced by the poor performance of the Amazon-Google and Walmart-Amazon models on the WDC dataset, as shown in Table 2. Considering that the larger models are only fine-tuned using the WDC dataset, in-domain generalization can only be assessed within the product domain. The fine-tuned Llama 70B model demonstrates the ability to generalize to the Amazon-Google and Walmart-Amazon datasets, as evidenced by its performance exceeding the baseline by 3.92 and 4.94 F1 points, respectively. However, it underperforms on the Abt-Buy dataset, with a decrease of 1.18 F1 points. Thus, fine-tuning improves performance on two out of three in-domain datasets.
In contrast, GPT-4o, while showcasing higher performance gains on the WDC dataset, exhibits poorer in-domain generalization. The fine-tuned version only manages to outperform the baseline on one out of three datasets.
@@
The Llama 8B models achieve 62% of the dedicated performance when generalizing within the scholar domain, while GPT-4o-mini attains 66%. These higher generalization scores are likely due to the higher similarity among the scholar datasets, as both benchmarks include records from DBLP.
@@
Cross-Domain Generalization: In a product-to-scholar cross-domain generalization scenario, Llama 8B models experience an average decrease of 11.05 F1 points, while GPT-4o-mini models show a decrease of 2.07 F1 points compared to their zero-shot performance. Specifically, all eight Llama 8B models and four out of eight GPT-4o-mini models underperform relative to their zero-shot baselines, as seen in the DBLP-ACM and DBLP-Scholar columns of Table 2.
@@
This trend continues with the scholar versions of the models. Llama 8B models underperform their zero-shot counterparts in the product domain by 1.11 F1 points, whereas GPT-4o-mini models gain an average of 1.60 F1 points over zero-shot. Larger models exhibit similar patterns: Llama 70B fails to surpass zero-shot performance in cross-domain scenarios, while GPT-4o gains an average of 4.12 F1 points, as reflected in the scholar domain columns of Table 2.
@@
In conclusion, standard fine-tuning significantly enhances in-domain generalization for the smaller models, while the results for the larger models as well as for cross-domain generalization are mixed.
3.3. Prompt Sensitivity
Variations of the wording of a prompt can have considerable performance implications on the model (Narayan et al., 2022b; Zhao et al., 2021; Peeters and Bizer, 2023; Liu et al., 2023). It is recommended to use the same prompt for fine-tuning that will later be used for querying the model  [^5] https://platform.openai.com/docs/guides/fine-tuning. In this section, we investigate how the performance of the fine-tuned models changes when a different prompt is used for querying the model.
Prompts: We assess the performance of the following three alternative prompts for querying models that were fine-tuned using the prompt shown in Figure 2:
  • simple-free: “Do the two product descriptions match?”
  • complex-force: “Do the two product descriptions refer to the same real-world product? Answer with ‘Yes’ if they do and ‘No’ if they do not.”
  • simple-force: “Do the two product descriptions match? Answer with ‘Yes’ if they do and ‘No’ if they do not.”
We measure the prompt sensitivity as the standard deviation of the performance across the prompt that is used for fine-tuning and the three alternative prompts. The detailed prompt sensitivity results can be found in the accompanying repository. Post-fine-tuning Llama 8B’s standard deviation drops from 15.76 to 1.87 in a non-transfer setting. The sensitivity increases slightly for in-domain transfer scenarios, with product models at 2.27 F1 points and scholar models at 2.60 F1 points. Fine-tuned Llama models show an average prompt sensitivity of 3.54 F1 points, down from 15.76 pre-fine-tuning. Furthermore, the prompt used for fine-tuning does not necessarily perform best with the alternative prompts resulting in the highest F1 score in 29 out of 42 scenarios for Llama.
GPT-4o-mini shows a similar trend, with a pre-fine-tuning prompt sensitivity of 2.72 F1 points. Post-fine-tuning, prompt sensitivity is reduced to 0.26 F1 points in non-transfer scenarios, 0.85 F1 points in the product domain, and 0.18 F1 points in the scholar domain. Across all datasets, fine-tuned GPT-4o-mini models exhibites a prompt sensitivity of 1.31 F1 points. With the prompt used for fine-tuning reporting the highest result in 50% of cases.
4. Dimension 1: Example Representation

### v2

@@ Abstract
Generative large language models (LLMs) are a promising alternative to pre-trained language models for entity matching due to their high zero-shot performance and ability to generalize to unseen entities. Existing research on using LLMs for entity matching has focused on prompt engineering and in-context learning. This paper explores the potential of fine-tuning LLMs for entity matching. We analyze fine-tuning along two dimensions: 1) the representation of training examples, where we experiment with adding different types of LLM-generated explanations to the training set, and 2) the selection and generation of training examples using LLMs. In addition to the matching performance on the source dataset, we investigate how fine-tuning affects the model’s ability to generalize to other in-domain datasets as well as across topical domains. Our experiments show that fine-tuning significantly improves the performance of the smaller models while the results for the larger models are mixed. Fine-tuning also improves the generalization to in-domain datasets while hurting cross-domain transfer. We show that adding structured explanations to the training set has a positive impact on the performance of three out of four LLMs, while the proposed example selection and generation methods, only improve the performance of Llama 3.1 8B while decreasing the performance of GPT-4o-mini.
Index Terms:
Entity matching, identity resolution, large language model, fine-tuning.
I Introduction
@@
Entity matching [1, 2, 3] is the task of identifying entity descriptions in different data sources that refer to the same real-world entity. Entity matching is a central step in data integration pipelines [4]. Many state-of-the-art entity matching methods build on pre-trained language models (PLMs) [5], such as BERT or RoBERTa [6, 7, 8, 9]. Recent work on using generative large language models (LLMs) [10], such as GPT, Llama, Gemini, or Mistral, for entity matching has shown that LLMs have higher zero-shot performance compared to PLMs and generalize better to unseen entities [11, 12, 13]. Most research on using LLMs for entity matching focuses on prompt engineering and in-context learning [11, 12, 14]. Initial papers have started to investigate the potential of improving beyond in-context learning by fine-tuning LLMs for entity matching [13, 15]. This paper builds on this work and presents a deeper investigation of the utility of fine-tuning LLMs for entity matching. Besides the effect of fine-tuning on the matching performance, we investigate how fine-tuning influences the model’s ability to generalize to different in-domain datasets as well as across topical domains. Our investigation focuses on two dimensions: First, the representation of training examples, where we investigate the effect of augmenting each training example with textual as well as structured explanations about why a pair of entity descriptions matches or not. The second dimension is the selection and generation of examples for the training set that is used for fine-tuning. We experiment with using LLMs to filter irrelevant examples from the training set as well as to generate additional training examples. We further investigate the error-based selection of training examples. Figure 1 gives an overview of the fine-tuning and inference setup. The effects of fine-tuning are compared between open-source (Llama-3.1) and proprietary (GPT-4o) models.
@@
Figure 1: Fine-tuning and inference setup including explanation generation, example generation, and filtration.
@@
  2. 2.
    In-domain and Cross-domain Generalization: We assess the effectiveness of fine-tuned models in both in-domain and cross-domain settings. The experiments show that fine-tuned models often generalize better to in-domain datasets, while their performance in cross-domain settings falls short of the zero-shot baseline in the majority of cases.
  3. 3.
@@
  4. 4.
    Example Selection and Generation: We introduce various approaches for using LLMs to filter the training set and generate new examples. While these methods yield notable gains for Llama 3.1 8B, they result in performance degradation for GPT-4o-mini.
II Experimental Setup
This section provides details on the models, benchmark datasets, fine-tuning setup, and the performance evaluation that we use for the experiments. All artifacts are provided in the project repository, meaning that all experiments can be replicated [^1] https://github.com/wbsg-uni-mannheim/TailorMatch.
@@ II Experimental Setup
  1. 1.
    gpt-4o-2024-08-06: This LLM was released in August 2024 as OpenAI’s flagship model and is used by ChatGPT.
  2. 2.
    gpt-4o-mini-2024-07-18: This hosted LLM, released in July 2024, has lower usage fees compared to GPT-4o. Although the model size is not disclosed, the naming convention indicates that it is smaller than GPT-4o.
  3. 3.
@@
  4. 4.
    Meta-Llama-3.1-70B-Instruct333meta-llama/Meta-Llama-3.1-70B-Instruct: Also developed by Meta, this model is medium-sized within the Llama series. Further details are available in the accompanying paper by Meta [16].
We use the OpenAI API [^4] https://platform.openai.com/docs/guides/batch and Hugging Face’s Transformers [^5] https://huggingface.co/docs/transformers/en/index library to prompt both hosted and local models. In both cases, the LLM temperature is set to zero in order to minimize randomness. We run open-source models on a multi-GPU server equipped with up to eight Nvidia RTX6000 GPUs, using one GPU for LLaMA 8B and four GPUs for the 70B model.
Datasets: We evaluate in-domain and cross-domain generalization using benchmark datasets from two topical domains [17, 8]: products and scholarly works. Each dataset has at least 150 matches in its test set (Table I) to ensure stable measurement. Second, we prioritize datasets that contain a significant number of challenging corner case record pairs. These corner cases are either matching or non-matching pairs that closely resemble the opposite class due to very similar or dissimilar surface forms [8]. Finally, to evaluate cross-domain generalization, we chose datasets from two topical domains. We use the following datasets:
@@
  • Walmart-Amazon: This benchmark includes products from both Walmart and Amazon, with categories comparable to those in WDC Products and Abt-Buy.
@@
TABLE I: Dataset statistics for the training, validation, and test sets.
@@
Fine-tuning Setup: We use the default hyperparameter setting recommended by the model providers. The fine-tuning setup differs for hosted and open-source models. We use the following default hyperparameter settings for the hosted OpenAI models: learning rate multiplier 1.8, batch size 16. We use the following fine-tuning setup for the open-source models: We employed Low-Rank Adaptation (LoRA) with an alpha of 16 to allow moderate adjustments to model weights, a dropout rate of 0.1 to prevent overfitting, and a rank (r) of 64 to balance performance and computational efficiency. The learning rate was configured at 2e-4. Both open-source and hosted models are trained for 10 epochs, with checkpoints generated after each epoch. Open-source models are validated at each checkpoint using custom evaluation callbacks. In contrast, OpenAI models provided only two intermediate checkpoints plus the final version, limiting the validation process.
Evaluation: Following related work [1], we use precision, recall, and F1. All subsequent tables show the F1 performance, while precision and recall results are provided in the accompanying repository. Model responses in natural language are evaluated using Narayan et al.’s [11] method of parsing responses for “yes” or “no” decisions.
III Standard Fine-Tuning
To establish baselines, we report results from fine-tuning experiments using a simple, straightforward representation of training examples. Figure 2 shows two examples from this set. Each example consists of a prompt asking if the entity descriptions refer to the same entity, followed by both descriptions and the completion “Yes” or “No” based on whether they match.
Figure 2: Representation of training examples for standard fine-tuning.
Table II reports results of standard fine-tuning experiments. Each row shows a specific LLM (column Model) fine-tuned on a given training set (column Training set). Subsequent columns display performance across multiple test sets. The zero-shot rows show results for models without fine-tuning.
TABLE II: F1 scores after standard fine-tuning. The rows determine the model/training set combination, while the columns contain results for different test sets. The best result of each model per test set is set bold, the second best result is underlined.
Model Training set Product Domain Average Scholar Domain Average
@@ III-A Effectiveness
Smaller Models: Llama-8B shows a 17.31-point average F1 gain over zero-shot baseline (Table II, upper section). 5 of 6 models significantly improve, with Abt-Buy achieving the highest gain (30.77 points, Table II, A-B column). Amazon-Google is an outlier with only 0.84 points improvement (Table II, A-G column). GPT-4o-mini demonstrates an 11.72-point average increase (Table II, rows gpt-4o-m). Unlike Llama-8B, GPT-4o-mini achieves substantial gains on Amazon-Google, its highest absolute improvement (Table II, A-G column).
@@
Large Models: Due to resource constraints, we fine-tune larger models only on the WDC small training set. Llama 70B experiences a performance decrease (-2.53 points, Table II, WDC column). In contrast, GPT-4o improves by 5.43 points to reach an F1 score of 87.07 on WDC, the highest result observed so far, despite starting with a similar zero-shot score as GPT-4o-mini.
III-B Generalization
@@
We evaluate two types of generalization: in-domain, meaning that generalization is assessed across datasets within the same topical domain, and cross-domain, where a model is trained using a training set from one domain and tested with a test set belonging to a dataset from a different topical domain (e.g., from products to scholar). In order to measure how well a model that was fine-tuned using training data from dataset ($D_{\text{source}}$) generalizes to a target dataset ($D_{\text{target}}$), we compare the performance of the model before and after fine-tuning (performance gain) to the performance gain of a model that was fine-tuned directly for $D_{\text{target}}$ using the train set of $D_{\text{target}}$. Let:
  • $F_{0}^{(t)}$ be the zero-shot F1 score of a model on $D_{\text{target}}$,
  • $F_{\text{transfer}}^{(t)}$ the F1 of a model fine-tuned on $D_{\text{source}}$ but evaluated on $D_{\text{target}}$, and
  • $F_{\text{target}}^{(t)}$ the F1 of a model fine-tuned and evaluated on $D_{\text{target}}$,
For each target dataset $t$, we define the transfer gain:
Transfer $\text{Transfer gain}(t)\;=\;\frac{F_{\text{transfer}}^{(t)}\;-\;F_{0}^{(t)}}{F_{\text{target}}^{(t)}\;-\;F_{0}^{(t)}}$ (1)
The transfer gain measures which fraction of the performance gain that can be achieved by fine-tuning using training data from $D_{\text{target}}$ is achieved by fine-tuning using training data from $D_{\text{source}}$. Since we have multiple target datasets from the same topical domain (e.g., product-related datasets like Abt-Buy and Walmart-Amazon), we report the average of all transfer gains from a source dataset $D_{\text{source}}$ to all target datasets ($T$) within a topical domain in the following.
In-domain Generalization: In the product domain, 5/6 Llama 8B models improve performance across both their own and other datasets. Amazon-Google is the only model showing decline in a in-domain generalization setting (Table II, WDC column). On average, Llama 8B models achieve 59% of the performance in an in-domain transfer scenario compared to models that were fine-tuned directly for the target datasets.
GPT-4o-mini exhibits less transfer capability, achieving on average a only 15% performance increase relative to dataset-specific models, primarily due to poor Amazon-Google and Walmart-Amazon performance on WDC (Table II).
For larger models (fine-tuned only on WDC), Llama 70B generalizes successfully to Amazon-Google and Walmart-Amazon (+3.92 and +4.94 F1) but underperforms on Abt-Buy (-1.18). GPT-4o, despite higher WDC gains, generalizes poorly, outperforming baseline on only 1/3 datasets.
@@
In the scholar domain, Llama 8B and GPT-4o-mini achieve better generalization (62% and 66% respectively), likely due to higher dataset similarity, as both benchmarks include DBLP records.
@@
Cross-Domain Generalization: In product-to-scholar transfer, Llama 8B models average a decrease of 11.05 F1 points, while GPT-4o-mini shows a 2.07-point decrease compared to zero-shot. All Llama 8B and half of GPT-4o-mini models underperform their zero-shot baselines (Table II, DBLP-ACM and DBLP-Scholar columns).
@@
In scholar-to-product transfer, Llama 8B models underperform by 1.11 F1 points, while GPT-4o-mini gains 1.60 points on average. Larger models follow similar patterns: Llama 70B fails to exceed zero-shot in cross-domain scenarios, while GPT-4o gains 4.12 F1 points (Table II, scholar domain columns).
@@
In conclusion, standard fine-tuning significantly enhances in-domain generalization for smaller models, while results for larger models and cross-domain generalization remain mixed.
IV Dimension 1: Example Representation